# Low-temperature thermal transport in moiré superlattices

### v1

@@ IV Conclusion
We have studied the thermal transport of various moiré superlattice systems in the low-temperature regime. In general, we found a reduction of thermal conductivity due to an overall flattening of the phonon bands. The largest reduction is around 35-40% which occurs at around 20 K for 0.817 TBG and 0 t-G/hBN and around 4 K for 0.8 t-MoS2/MoS2. At higher temperature, the thermal conductivity returns towards the original intrinsic value, as the moiré superlattice effect for the acoustic phonons is significant only in the low energy region. These changes result in a characteristic deviation to the original quadratic temperature dependence of two-dimensional systems where linear acoustic phonons dominate the thermal transport properties. The results could provide a pathway for engineering thermal conductivity and observation of the moiré effect in the phonon properties of twisted bilayer system. While the present study is limited in the harmonic approximation, we expect that anharmonic scattering between moiré phonons and the presence of disorder [70, 72] could further decrease the thermal conductivity. We leave this imposition for future study.

### v2

@@ IV Conclusion
We have studied the phonon thermal conductivity of various moiré superlattice systems in the low-temperature regime.
Using the continuum model, we showed that half of the in-plane phonon modes are hybridized by the interlayer moiré potential with notable reduction in their group velocity, particularly at small twist angles. This leads to the decrease of thermal conductivity across a wide temperature range. For example, in the case of 0 t-G/hBN, the highest reduction occurs at around 20 K, for up to 40%, and this reduction slowly subsides at higher temperatures until it returns to the original intrinsic value at around 100 K. On the contrary, at very low temperatures, thermal conductivity is shown to be inversely related to the phonon velocity, and therefore it becomes enhanced. This continuous change over temperature leads to a power-law deviation from the original quadratic temperature dependence, that is, the exponent $\alpha$ is less than 2 from 4 K to 20 K and greater than 2 from 20 K to 80 K in the case of 0 t-G/hBN. This behavior also holds in other moiré systems, though the temperature ranges in which it appears depend on the original acoustic phonon velocity of the constituent monolayers.
We expect future observation of the characteristic power-law deviation to be strong evidence of moiré effect on low-energy phonons, further highlighting its importance when investigating phonon-related phenomena in moiré systems. Thermal conductivity measurements at low temperatures down to a few kelvins have been demonstrated, for example, in carbon nanotubes [72] and thin graphite [73]. While distinguishing the contribution of in-plane and out-of-plane modes in thermal transport experiments is generally challenging, one could consider cases where the mean free path of the in-plane modes is significantly larger, such as in a $\mu$m-wide sample supported or sandwiched by substrates [38, 74, 42, 43, 44].
Lastly, the inclusion of anharmonic scattering between moiré phonons and the presence of moiré-scale disorder [70, 75] is also expected to further decrease the thermal conductivity. We leave this investigation for future study.